OpenStudy (anonymous):

If z varies jointly with x and y and inversely with w, and z = 7 when x = 2, y= -1, and w = 4, find z when x =3, y = 4, and w =6

OpenStudy (nikato):

WELCOME TO OPENSTUDY! so z varies jointly with x and y and inversely with w is \[z=\frac{ kxy }{ w }\]

OpenStudy (nikato):

and you're given z=7,x=2,and w=4,y=-1 now plug these values in and can you find k? \[7=\frac{ k(2)(-1) }{ 4 }\]

OpenStudy (anonymous):

Is it -14?

OpenStudy (johnweldon1993):

Indeed ^ k = -14 So now you have \[\large z = \frac{-14xy}{w}\] Plug in the value you have and the values you want...to solve for 'z' x = 3 y = 4 w = 6 \[\large z = \frac{-14(3)(4)}{6}\]

OpenStudy (nikato):

correct so now your equation is \[z=\frac{ -14xy }{ w }\] now use this formula by pluggin in your known values to solve for z

OpenStudy (anonymous):

Ah thank you guys! ;v; I know how to solve these now

OpenStudy (nikato):

you're welcome!